Dear Tutor, may you please help me on this problem:
The wheels of a bicycle made 50 revolutions to travel a distance of 88 m. If the wheels of this bicycle are identical, what is the diameter of the wheel?
Distance covered in one revolution=_____________m=______m
Diameter of the wheel =______________m= _______m
Thank you!

Answer by lwsshak3(11628)

The wheels of a bicycle made 50 revolutions to travel a distance of 88 m. If the wheels of this bicycle are identical, what is the diameter of the wheel?
**
one revolution=circumference of bicycle wheel=π*diameter
50 revolutions=50 circumferences=50*π*diameter=88 m
diameter=88/50π≈.560 m
Distance covered by one revolution=π*.560 m